OpenStudy (anonymous):

The table below shows two equations: Equation 1 |4x – 3|− 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = -4, 1. The solutions to equation 1 are x = 3, -1.5 and equation 2 has no solution. The solutions to equation 1 are x = 3, -1.5 and equation 2 has solutions x = -4, 1. please help in this one please

OpenStudy (cwrw238):

hint : the modulus functions |4x – 3| and |2x + 3| by definition are positive

OpenStudy (cwrw238):

can u solve equation 1?

OpenStudy (anonymous):

i really dont know this one please can you just give me the correct letter?

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

to solve equation 1 4x - 3 = 9 or 4x - 3 = -9 equation 2 we have a positive value + 8 = 3 is this possible?

OpenStudy (anonymous):

Equation 1 |4x – 3|− 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = -4, 1. The solutions to equation 1 are x = 3, -1.5 and equation 2 has no solution. The solutions to equation 1 are x = 3, -1.5 and equation 2 has solutions x = -4, 1.

OpenStudy (anonymous):

tell me wich letter

OpenStudy (cwrw238):

the purpose of this site is to help you how to solve problems not just give you the answer

OpenStudy (anonymous):

please ? i need to pass this test ?

OpenStudy (cwrw238):

i have helped you a lot equation 1 can be solved and i've given you a good clue about equation 2

OpenStudy (tjhub00):

the answer is c